# HARP: Hesitation-Aware Reframing in Transformer Inference Pass

### v1

@@ Abstract
This paper aims to improve the performance of large language models by addressing the variable computational demands in inference steps, where some tokens require more computational resources than others. We present HARP, a simple modification to “off-the-shelf” Transformer forward pass. Drawing from hesitation and the framing effect in decision-making, HARP selectively applies additional computation when the model encounters uncertainty during token generation. Our method mimics human cognitive processes by pausing at difficult decision points and reframing inputs for a different perspective. Unlike other approaches, HARP is model-agnostic, training-free, and easy to implement. We thoroughly evaluate our method across various downstream tasks and model sizes, demonstrating performance improvements up to +5.16%. Notably, HARP achieves these gains while maintaining inference times twice faster than beam search. Simple and yet with significant gains, HARP offers a practical solution for enhancing the performance of Transformer-based language models with minimal computational impact.
@@ 1 Introduction
  • We combine these two components to create HARP (Hesitation-Aware Reframed Forward Pass), a method that generally applies to any decoding algorithms, and thoroughly evaluate its performance across various tasks, model sizes, and in combination with existing techniques.
@@ 3.4 Hesitation-Aware Reframed Forward Pass (HARP)
However, if the uncertainty exceeds the threshold $\theta$, the model is uncertain. In this case, we initiate a second forward pass but with a reframed input sequence (explained in Section 3.3). The reframed embeddings, $\mathbf{\hat{e}}$, are obtained by applying dropout with a rate $\delta$ to the original embeddings $\mathbf{e}$. We then perform a second forward pass using $\mathbf{\hat{e}}$, which outputs reframed logits, denoted as $logits_{r}$.
@@
Our method introduces three hyperparameters: the uncertainty threshold $\theta$, the dropout rate $\delta$, and the combination factor $\beta$. In our experiments, we set these parameters to values identified as optimal during preliminary testing. The optimal choices for these hyperparameters may vary across tasks or models and can benefit from further tuning; we leave studying the impact of these values in more depth as future work.
@@ 4.2 Datasets
Prompt details can be found in Appendix C. All the datasets are evaluated with a zero-shot prompt.
@@ 4.3 Evaluation
The hyperparameters used in our experiments are fixed at a dropout rate of $\delta=0.20$ and an uncertainty threshold of $\theta=1.0$. We set the temperature for nucleus sampling to $\tau=0.6$ and the top-$p$ to 0.9. In the case of beam search, the number of beams is $b=3$ with a top-$k$ of 5. We chose to apply length-normalization (Wu et al., 2016) as we found it beneficial to tasks such as summarization and reasoning. This length penalty is fixed to $\alpha=0.6$ for every datasets, as we seek to compare HARP with model- and task-agnostic methods. We empirically found the value by evaluating a model on a subset of three datasets (CNN/DM, GSM8K and LAMBADA). When evaluating Chain-of-Thought (Wei et al., 2024) prompting, we prepend “Let’s think step-by-step.” to the generation of the model. We evaluate a subset of each dataset with a batch size of 1, without caching, on a single RTX3090 (24GB) using the same seed for every example. The accuracy is reported for all datasets except CNN/DM, where we evaluate the ROUGE-1 score. Moreover, we record the generation time of each method.
5 Results
@@
In summary, our HARP allows further improvements as high as +5.16% of high-ended models that already show strong performance on many tasks without requiring any fine-tuning. This demonstrates the potential of HARP to enhance state-of-the-art models.
@@
As shown in Table 2, advanced techniques, such as Chain-of-Thought (CoT) (Wei et al., 2024) prompting, are further enhanced with HARP. For instance, applying the modified forward pass to CoT prompting with the LLaMA model improves accuracy by 0.79%, 1.16%, and 1.09% on CommonsenseQA, GSM8K, and MMLU Pro, respectively. This shows our method can seamlessly combine with other existing methods, further enhancing model performance.
@@
Figure 2 reveals that despite the additional computation, the inference time of HARP remains close to that of the vanilla generation, introducing a minimal additional computational cost. The inference time for our method is consistently under twice that of the original model across all tasks. In contrast, beam search—while most of the time yielding lower performance than HARP—induces a significantly higher inference cost. This is the indicator that our method strikes a balance between performance gains and inference time.
Beam search shows significant delays, often higher than x2.4 the time required by the greedy search generation, whereas HARP achieves inference times lower than x1.4. Additional results in Table 6 validate this efficiency by comparing other models. For example, the Mistral model increases inference time by only x1.17 to x1.49, while beam search increases it by x2.51 to x3.72.
@@ 6 Analysis
Although GSM8K achieves a notable +2.00% increase in accuracy using four additional steps with LLaMA, these gains come at the cost of significantly higher computation time, scaling up to x1.97 in the 4-step method.
We hypothesize that more than two representations might confuse and distract the model from the original task, decreasing precision. This aligns with cognitive theories, suggesting that while considering reframings enhances problem-solving, too many representations increase cognitive load (Sweller, 1988), reducing performance. In a more mathematical way, this behavior may be due to the random noise introduced during reframing and the way logits are combined. While random noise can sometimes be beneficial, it can also be detrimental, and increasing the number of reframings increases the likelihood of harmful noise. Additionally, as the logits from the vanilla and reframed representations are averaged (see Equation 4), introducing more than one reframing step reduces the weight of the vanilla logits, disproportionately favoring the reframed ones, contributing the the higher likelihood of harmful representations.
@@ 7 Conclusion
This paper presented a novel method, HARP, designed to enhance language model inference without requiring retraining or architectural modifications. Our approach uses a selection process based on token-level uncertainty to identify when additional computation is advantageous and an innovative method for reframing inputs during inference. We demonstrated that HARP significantly improves performance across various tasks and model sizes by combining these components, achieving accuracy gains up to 5.16%. Importantly, HARP maintains inference efficiency compared to methods like beam search, with only an x1.25 average increase in inference time over the standard model.
@@ 8 Limitations
First, due to resource and time constraints, the evaluation of HARP was conducted on a limited subset of each dataset. Additionally, the method has yet to be tested on larger language models, particularly those with 70 billion parameters or more. Although we attempt to cover a range of tasks, there are other language challenges that our method has not been tested on. Consequently, we cannot definitively assess how HARP generalizes across broader tasks or scales to significantly larger models.
@@ Appendix B Uncertainty in practice
Appendix C Prompts

### v2

@@ Abstract
This paper aims to improve the performance of large language models by addressing the variable computational demands in inference steps, where some tokens require more computational resources than others. We present HARP, a simple modification to “off-the-shelf” Transformer forward pass. Drawing from hesitation and the framing effect in decision-making, HARP selectively applies additional computation when the model encounters uncertainty during token generation. Our method mimics human cognitive processes by pausing at difficult decision points and reframing inputs for a different perspective. Unlike other approaches, HARP is model-agnostic, training-free, and easy to implement. We evaluate our method across various downstream tasks and model sizes, demonstrating performance improvements up to +5.16%. Notably, HARP achieves these gains while maintaining inference times twice faster than beam search. Simple and yet with significant gains, HARP provides insights into the potential of adaptive computation for enhancing the performance of Transformer-based language models.
@@ 1 Introduction
  • We combine these two components to create HARP (Hesitation-Aware Reframed Forward Pass), a method that generally applies to any decoding algorithms, and evaluate its performance across various tasks, model sizes, and in combination with existing techniques.
@@ 3.4 Hesitation-Aware Reframed Forward Pass (HARP)
However, if the uncertainty exceeds the threshold $\theta$, the model is uncertain. In this case, we initiate a second forward pass but with a reframed input sequence (explained in Section 3.3). The reframed embeddings, $\mathbf{\hat{e}}$, are obtained by applying dropout with a rate $\delta$ to the original embeddings $\mathbf{e}$. We then perform a second forward pass using $\mathbf{\hat{e}}$, which outputs reframed logits, denoted as $logits_{r}$. This additional forward pass requires recomputation as it cannot leverage the KVCache (see Section 8).
@@
Our method introduces three hyperparameters: the uncertainty threshold $\theta$, the dropout rate $\delta$, and the combination factor $\beta$. In our experiments, we set these parameters to values identified as optimal during preliminary testing. The optimal choices for these hyperparameters may vary across tasks or models and can benefit from further tuning; we provide a preliminary study on the impact of $\theta$ in Appendix C and leave a more in-depth exploration of all hyperparameters for future work.
@@ 4.2 Datasets
Prompt details can be found in Appendix E. All the datasets are evaluated with a zero-shot prompt.
@@ 4.3 Evaluation
The hyperparameters used in our experiments are fixed at a dropout rate of $\delta=0.20$ and an uncertainty threshold of $\theta=1.0$. We set the temperature for nucleus sampling to $\tau=0.6$ and the top-$p$ to 0.9. In the case of beam search, the number of beams is $b=3$ with a top-$k$ of 5. We chose to apply length-normalization (Wu et al., 2016) as we found it beneficial to tasks such as summarization and reasoning. This length penalty is fixed to $\alpha=0.6$ for every datasets, as we seek to compare HARP with model- and task-agnostic methods. We empirically found the value by evaluating a model on a subset of three datasets (CNN/DM, GSM8K and LAMBADA). When evaluating Chain-of-Thought (Wei et al., 2024) prompting, we prepend “Let’s think step-by-step.” to the generation of the model. We evaluate a subset of each dataset with a batch size of 1, without caching (KVCache), on a single RTX3090 (24GB) using the same seed for every example. The accuracy is reported for all datasets except CNN/DM, where we evaluate the ROUGE-1 score. Moreover, we record the generation time of each method.
5 Results
@@
In summary, our HARP allows further improvements as high as +5.16% of high-end models that already show strong performance on many tasks without requiring any fine-tuning. This demonstrates the potential of HARP to enhance state-of-the-art models.
@@
As shown in Table 2, advanced techniques, such as Chain-of-Thought (CoT) (Wei et al., 2024) prompting, are further enhanced with HARP. For instance, applying the modified forward pass to CoT prompting with the LLaMA model improves accuracy by 0.8%, 1.16%, and 1.10% on CommonsenseQA, MMLU Pro, and GSM8K, respectively. This shows our method can seamlessly combine with other existing methods, further enhancing performance.
@@
Figure 2 reveals that despite the additional computation, the inference time of HARP remains close to that of the vanilla generation, introducing a minimal additional computational cost. The inference time for our method is consistently under twice that of the original model without caching across all tasks. In contrast, beam search—while most of the time yielding lower performance than HARP—induces a significantly higher inference cost. This is the indicator that our method strikes a balance between performance gains and inference time.
Beam search shows significant delays, often higher than x2.4 the time required by the greedy search generation, whereas HARP achieves inference times lower than x1.4. Additional results in Table 6 validate this efficiency by comparing other models. For example, the Mistral model increases inference time by only x1.17 to x1.49, while beam search increases it by x2.51 to x3.72, compared to the original model without KVCache.
@@ 6 Analysis
Although GSM8K achieves a notable +2.01% increase in accuracy using four additional steps with LLaMA, these gains come at the cost of significantly higher computation time, scaling up to x1.97 in the 4-step method.
We hypothesize that more than two representations might confuse and distract the model from the original task, decreasing precision. This aligns with cognitive theories, suggesting that while considering reframings enhances problem-solving, too many representations increase cognitive load (Sweller, 1988), reducing performance. In a more mathematical way, this behavior may be due to the random noise introduced during reframing and the way logits are combined. While random noise can sometimes be beneficial, it can also be detrimental, and increasing the number of reframings increases the likelihood of harmful noise. Additionally, as the logits from the vanilla and reframed representations are averaged (see Equation 4), introducing more than one reframing step reduces the weight of the vanilla logits, disproportionately favoring the reframed ones, contributing to the higher likelihood of harmful representations.
@@ 7 Conclusion
This paper presented a novel method, HARP, designed to enhance language model inference without requiring retraining or architectural modifications. Our approach uses a selection process based on token-level uncertainty to identify when additional computation is advantageous and an innovative method for reframing inputs during inference. We demonstrated that HARP can achieve significant performance improvements across various tasks and model sizes, with accuracy gains of up to 5.16%. Importantly, HARP maintains inference efficiency compared to methods like beam search, with only an x1.25 average increase in inference time over the standard model. While HARP provides a promising proof of concept, its real-world application is currently limited by challenges such as cache invalidity and the introduced randomness.
@@ 8 Limitations
First, due to resource and time constraints, our experiments were limited in scope. Evaluations were conducted on quantized models and subsets of each dataset, and we implemented custom generation methods. Additionally, the method has yet to be tested on larger language models, particularly those with 70 billion parameters or more. Although we attempt to cover a range of tasks, there are other language challenges that our method has not been tested on. Furthermore, our experiments did not leverage widely used libraries, such as LM-Evaluation-Harness (Gao et al., 2021), which could provide more standardized benchmarking. Consequently, while initial results demonstrate potential, further work is required to confirm whether HARP generalizes effectively across tasks and scales. To partially address this limitation, we provide an extended evaluation using LM-Evaluation-Harness in Appendix D. HARP serves as a proof of concept, demonstrating the potential of uncertainty-aware adaptive computation in improving inference performance.
@@ Appendix B Uncertainty in practice
Appendix C Impact of the hesitation threshold
Datasets Value of $\theta$
0.6 0.8 1.0 1.2 1.4 1.6 1.8 2.0
CsQA x1.03 x1.02 x1.01 x1.00 x1.00 x1.00 x1.00 x1.00
LAMBADA x1.53 x1.49 x1.40 x1.33 x1.31 x1.29 x1.27 x1.26
Table 7: Relative inference time of HARP using LLaMA-3.1 Instruct 8B with different values of $\theta$ compared to the Vanilla method.
Datasets Value of $\theta$
0.6 0.8 1.0 1.2 1.4 1.6 1.8 2.0
CsQA +1.72 +2.12 +1.52 +1.31 +1.72 +1.52 +1.41 +1.31
LAMBADA +5.12 +5.21 +5.16 +5.12 +5.08 +5.02 +4.95 +4.85
Table 8: Relative accuracy of HARP using LLaMA-3.1 Instruct 8B with different values of $\theta$ compared to the Vanilla method.
Tuning the hyperparameter $\theta$ is beyond the scope of this paper. However, we provide a preliminary study to analyze its impact. We evaluate two datasets with different $\theta$ values and report relative accuracy and inference time in Table 8 and Table 8 respectively.
As $\theta$ increases, inference speed improves thanks to fewer reframing steps—leading to reduced computation. However, this comes at the cost of lower accuracy. Conversely, a too small $\theta$ can also degrade accuracy. This aligns with the discussion in Section 6 on multi-step reframing, where additional reframing may cause the model to “overthink” and lose track of improvements.
This preliminary study is based on a limited set of tasks, and further investigation is needed to better understand the selection of $\theta$.
Appendix D Extended Evaluation
To further validate the robustness of our method, we conducted additional evaluations of HARP using the LM-Evaluation-Harness (Gao et al., 2021). This complements the original results in Section 4 by offering a broader and more standardized assessment.
All experiments were conducted on the full datasets using a single LLaMA-3.1 8B Instruct model (Dubey et al., 2024) in FP16 precision, without quantization, on an RTX A6000 GPU. Each evaluation used six different random seeds to assess variability, except for MMLU Pro, which was run with one seed due to its computational cost.
CommonsenseQA and LAMBADA were evaluated using generation-based decoding, consistent with our initial setup, rather than LM-Eval’s default log-likelihood evaluation. We report mean performance and standard deviation in Table 9.
Method CsQA GSM8K LAMBADA MMLU Pro CNN/DM
Vanilla (Greedy) 75.92 78.39 55.52 33.30 35.77
Ours (Greedy) 76.19 ($\pm$ 0.39) 78.57 ($\pm$ 0.66) 56.62 ($\pm$ 0.26) 33.56 36.13 ($\pm$ 0.03)
Table 9: Extended evaluation results using LM-Evaluation-Harness. Scores reflect Exact Match for CommonsenseQA, MMLU Pro, and LAMBADA; accuracy for GSM8K; and ROUGE-1 for CNN/DM.
These results confirm that HARP brings consistent gains across tasks. However, they also highlight a key limitation: the introduction of randomness. In particular, even when using greedy decoding, dropout-based reframing introduces stochasticity. While the average performance improves, certain seeds can still negatively impact results. This underscores the need for future work on reducing variance and improving robustness.
Appendix E Prompts